Subject: Minutes of March 9, 1998 APO User's Committee meeting

From: strauss@astro.Princeton.EDU

Submitted: Tue, 10 Mar 1998 10:33:20 -0500 (EST)

Message number: 260 (previous: 259, next: 261 up: Index)

	Apache Point Observatory 3.5m User's Committee Meeting
****************March 9, 1998**********************

Attending: Rene Walterbos, Michael Strauss, Bruce Gillespie, Ed
Turner, Jeff Secker, Ed Kibblewhite, Chris Stubbs

Not attending: Alan Uomoto

Gillespie: Report on a large number of fixes at APO. 

  The last ten days have been plagued by a number of problems, now
essentially under control. 
  Lowenstein has been working on Remark software.  Version 2.39 has
been released. 
  GRIM dithering scripts used to cause crashes of GRIM controlling
	computer; the problem seems to have been fixed. 
      Lucinio has done some work on the DIS, and improved some of the 
software for motor controllers.  One effect is that the nitrogen
consumption is down by a factor of 3! 
      Lowenstein is still working on putting medium-dispersion
gratings into Remark, but they are available to use now.
      Gloria has made measurements of DIS read noise: down by 30% on
red chip relative to values from February.  Now it is 14 electrons
(which is much better than it was several months ago).  No-one knows
how it got better!  One still sees evidence for banding in the bias
frames.  So there is still some sort of flakiness going on in the
analogue side.  What this really needs is an afternoon or two of Gunn's time.
      SPICAM shutter, which binds in various nasty ways, has been
fixed, at least as a kludge.  Stubbs has machined a new shutter blade
which can be used to replace it.  An alternative is to use an iris
shutter (which will give a small non-uniformity in the flat-field).
But in any case, it seems to be working now.
      The terminal server which controls the various motors of the
telescope, has given up the ghost (lost its memory).  A spare has been
put in, and seems to be working. 
      The telescope altitude drive died a few nights ago.  The APO
staff was able to get things back on line by the following evening.
There will be some expenses to re-obtain spares for these various things. 
      Network performance has gotten a bit better in the last few
months.  One of the bottlenecks, at Las Cruces, was gotten around by
using a different T1 line out of there.  That has improved things. 
      Engineering tests: drift-scanning code needs some serious
looking at.  But people have been very happy with the quality of
drift-scanning data; its nicely flat, and goes very deep. 
      It would be useful if one could drift-scan slower, to go
deeper.  It would also be useful to have a facility to figure out how
to drift between two pre-determined points on the sky. 
      The code can now drift-scan along great circles. 
      Latest pointing model: 2.5 arcsecond rms residuals. 

  At the moment, everything seems to be working great.  Let's hope
this continues for a while.

  Ae e-mail report from Uomoto given by Ed Turner: things are moving ahead on the secondary.
Mounting cell is more or less on schedule.  Initial payment has been
made to Steward, and contracts has been signed.  There has been some
slippage, but it is minor; we're still safely ensconced in the Steward

  Echelle: Current schedule: it will leave Chicago in May; working on
the telescope by the Summer.  There needs to be some real thinking
about the impact on the site of the Echelle commissioning effort.
Kibblewhite: Chicago and APO people should sit down together to make
sure everyone knows what is needed and what is available.  When would
be a good time to send some Chicago people down to talk to the APO
people?  Gillespie:  How about early May? 
  Gillespie: what else will come with the instrument: Software?
Documentation? User's manual? Spare parts? Remark modifications?
  Kibblewhite: Chicago folks will figure out a coherent answer to all
these concerns, and Doug Duncan or someone else will come down to APO
to discuss with the site staff.

  Stubbs: NSF instrumentation program proposal has deadline in
September or so.  Should we think about putting a proposal in for CCD
upgrades to Echelle, DIS?  We might want to pursue this, using UW
expertise of people working on Stubbs' wide-field camera, especially
Peter Dougherty.  One concern: It may be too early to apply to
upgrades to Echelle, before we can confirm that it is working in its
present form.  So let's wait a year for that, and for now, just put in
a proposal for DIS chip upgrades. 

There are plans for maintenance shutdown this summer, 3 weeks in July,
for work on telescope drives.  It's been 2 years since there has been
a substantial amount of work on GRIM.  Some of the things that people
would like: multiple-read capability, add some new filters, put in a
coronographic finger, general software update, so it might be good to
get Mark Hereld to the site during the shutdown to work on all of

  We will generate a policy page on the Web: how time is allocated and
charged to the different institutions, rules for when you can observe
remotely, etc. 

  Secker: Dome flats with SPICAM show large gradients: they look very
bad, presumably due to poor baffling. It would be useful to share
twilight flats between users.  Even sky flats are rotator angle and
elevation-dependent, because of lack of baffling.  Drift-scanning
makes it a 1-D problem, but it is still a problem.  A Nasymth conical
baffle is being built at NMSU.  

  There will be a phone con tomorrow morning to discuss current status
and next step of the 3-yr plan, led by Stubbs. 

  Stubbs: wide-field camera is something like 12-18 months away from
delivery, depending on the lead-time of the optics.  It will consist
of 3 rows of 8 detectors each.  There are options to make it fully
drift-scannable, or only having two of the three rows scannable (it is
not clear what the advantages if any, other than decreased complexity,
of the latter are); they would like community input on how important
drift-scanning with this instrument would be. 

  Previous month's minutes are approved. 

  Next meeting, April 13, 1998, 12:30 PM
APO APO APO APO APO  Apache Point Observatory 3.5m  APO APO APO
APO  This is message 260 in the apo35-general archive. You can find
APO  the archive on
APO  To join/leave the list, send mail to
APO  To post a message, mail it to